2) If we assume the π-network in decatetraene (C10H12) to serve as a one dimensional box containing eight π-electrons that can each be assumed to behave according to the particle in a box model where each level is assumed to be doubly degenerate, and if we further assume that the molecule is linear and use the values 135 and 154 pm for the C=C and C- C bonds respectively, what wavelength of light is required to induce a transition from the ground state to the first excited state?
Answer #1

CH CH = CH-C; CH-CH=CH-CH=CH-CH3 Decatetiocne e r dength of 5๙ bond 21 2 ーク の ro 8 mL- -31 22, - 31 1 0 x 1012 m)